# YOLO-based Bearing Fault Diagnosis With Continuous Wavelet Transform

### v1

@@ Abstract
This letter proposes a YOLO-based framework for spatial bearing fault diagnosis using time–frequency spectrograms derived from continuous wavelet transform (CWT). One-dimensional vibration signals are first transformed into time–frequency spectrograms using Morlet wavelets to capture transient fault signatures. These spectrograms are then processed by YOLOv9, v10, and v11 models to classify fault types. Evaluated on three benchmark datasets, including Case Western Reserve University (CWRU), Paderborn University (PU), and Intelligent Maintenance System (IMS), the proposed CWT–YOLO pipeline achieves significantly higher accuracy and generalizability than the baseline MCNN–LSTM model. Notably, YOLOv11 reaches mAP scores of 99.4% (CWRU), 97.8% (PU), and 99.5% (IMS). In addition, its region-aware detection mechanism enables direct visualization of fault locations in spectrograms, offering a practical solution for condition monitoring in rotating machinery.
@@
Bearing fault diagnosis, continuous wavelet transform (CWT), time–frequency spectrogram, YOLO object detection.
@@ I Introduction
Several traditional signal processing methods, such as short-time Fourier transform (STFT), Wigner–Ville distribution (WVD) [3], and Hilbert–Huang transform (HHT) [4], have been widely used to characterize non-stationary signals in both time and frequency domains. However, each of these methods exhibits intrinsic limitations [1, 2]. STFT employs a fixed analysis window, resulting in a fundamental trade-off between time and frequency resolution, which restricts its ability to capture features across multiple scales. WVD offers high resolution in theory, but it suffers from severe cross-term interference when analyzing multi-component signals. This interference often leads to spurious artifacts in the resulting spectrograms, which complicates the interpretation of fault features. HHT is data-adaptive and can decompose signals into intrinsic mode functions (IMFs), but its performance is highly sensitive to noise. Furthermore, it lacks a rigorous mathematical foundation for its empirical mode decomposition (EMD), which undermines the stability and reliability of its instantaneous frequency representation.
@@
To address these limitations, we propose a framework that combines CWT-based time-frequency spectrograms with You Only Look Once (YOLO)-based object detection models, including YOLOv9 [11], YOLOv10 [12], and YOLOv11 [13] to localize and classify fault patterns simultaneously. This approach leverages YOLO’s unified detection pipeline, which is optimized for both accuracy and speed, to enables fast and spatially explainable bearing fault diagnostics. To the best of our knowledge, this is the first study to explore YOLO-based spatial detection in time–frequency spectrograms for industrial bearing fault applications. The main contributions of this letter are as follows:
  • We propose a novel bearing fault diagnosis framework that combines CWT with YOLOv9, v10, and v11 to detect and classify fault patterns directly on time–frequency spectrograms.
@@ II-A Continuous Wavelet Transform (CWT)
In this work, we adopt the Morlet wavelet due to its favorable time–frequency localization and robustness to high-frequency noise. The Morlet wavelet combines a sinusoidal carrier with a Gaussian envelope $\psi(t)=e^{-\frac{t^{2}}{2}}\cos(5t)$, which offers balanced resolution across different frequency bands and has shown superior performance in early fault detection [6]. By applying CWT to raw vibration signals, we generate two-dimensional time–frequency spectrograms that expose transient features and scale-dependent energy distributions. These spectrograms serve as spatially structured inputs for subsequent object detection models.
II-B Time–Frequency Spectrogram Visualization
To validate the visual enhancement provided by CWT, we convert the same vibration signals from Fig. 1 into two-dimensional time–frequency spectrograms using Morlet-based wavelet transformation. The resulting images, shown in Fig. 2, capture distinct energy concentration patterns that correspond to different fault types.
@@ II-C1 Signal Segmentation and Spectrogram Generation
For each dataset (CWRU, PU, and IMS), the raw vibration signals are segmented into fixed-length windows, each containing a sufficient number of samples to capture fault-relevant transients. Empirically, a window length of $2048$ samples with $50\%$ overlap is used to balance time resolution and data diversity. Each segment is transformed into a 2D time–frequency spectrogram using the Morlet-based CWT described earlier. The resulting grayscale spectrograms are logarithmically compressed and normalized to a $[0,1]$ range.
@@ II-C2 Bounding Box Annotation for Object Detection
Unlike traditional classification models that produce a single label for an entire signal segment, YOLO models require localized object annotations in the form of bounding boxes. To enable this, we manually annotate each spectrogram using the LabelImg tool, enclosing the time–frequency regions that exhibit fault-specific energy concentrations. Each box is associated with a class label corresponding to one of four fault types: Normal, Ball fault, Inner race fault, and Outer race fault.
@@ III-C Performance Comparison Across Datasets
Compared to YOLOv9, YOLOv11 reduces FLOPs by over 97% and parameters by nearly 95%, while achieving comparable or better accuracy across all datasets. Although MCNN-LSTM remains the most compact model in terms of complexity, its classification accuracy and generalizability are significantly lower. In contrast, YOLOv11 provides a balanced trade-off between efficiency and high detection accuracy, making it suitable for lightweight industrial fault monitoring. The proposed CWT-YOLO framework demonstrates strong generalization across diverse datasets. On the PU dataset, which includes multiple fault types under varying load and speed, YOLOv11 outperforms the baseline by over 16 percentage points in mAP. On the IMS dataset, all YOLO models achieve near-perfect classification, highlighting their ability to capture subtle degradation over long-term operation. Among the YOLO variants, YOLOv11 achieves the best overall performance, combining lower model complexity with faster inference than YOLOv9 and YOLOv10. These results indicate that the spatial detection paradigm, when applied to time–frequency spectrograms, supports accurate and interpretable fault localization.
@@ IV Conclusion
This letter presents a bearing fault diagnosis framework that combines CWT with YOLO-based object detection to leverage time–frequency spectrograms for localized classification. By converting one-dimensional vibration signals into spatially structured spectrograms, the system enables visual identification of fault patterns using bounding box detection. Evaluations on the CWRU, PU, and IMS datasets confirm that YOLO-based models significantly outperform the MCNN-LSTM baseline in classification accuracy, with YOLOv11 achieving the best balance between detection performance and computational complexity. These results highlight the potential of spectrogram-level object detection for building practical, interpretable, and deployment-ready condition systems.

### v2

@@ Abstract
This letter proposes a YOLO-based framework for spatial bearing fault diagnosis using time-frequency spectrograms derived from continuous wavelet transform (CWT). One-dimensional vibration signals are first transformed into time-frequency spectrograms using Morlet wavelets to capture transient fault signatures. These spectrograms are then processed by YOLOv9, v10, and v11 models to classify fault types. Evaluated on three benchmark datasets, including Case Western Reserve University (CWRU), Paderborn University (PU), and Intelligent Maintenance System (IMS), the proposed CWT-YOLO pipeline achieves significantly higher accuracy and generalizability than the baseline MCNN-LSTM model. Notably, YOLOv11 reaches mAP scores of 99.4% (CWRU), 97.8% (PU), and 99.5% (IMS). In addition, its region-aware detection mechanism enables direct visualization of fault locations in spectrograms, offering a practical solution for condition monitoring in rotating machinery.
@@
Bearing fault diagnosis, continuous wavelet transform (CWT), time-frequency spectrogram, YOLO object detection.
@@ I Introduction
Several traditional signal processing methods, such as short-time Fourier transform (STFT), Wigner-Ville distribution (WVD) [4], and Hilbert-Huang transform (HHT) [5], have been widely used to characterize non-stationary signals in both time and frequency domains. However, each of these methods exhibits intrinsic limitations [1, 2]. STFT employs a fixed analysis window, resulting in a fundamental trade-off between time and frequency resolution, which restricts its ability to capture features across multiple scales. WVD offers high resolution in theory, but it suffers from severe cross-term interference when analyzing multi-component signals. This interference often leads to spurious artifacts in the resulting spectrograms, which complicates the interpretation of fault features. HHT is data-adaptive and can decompose signals into intrinsic mode functions (IMFs), but its performance is highly sensitive to noise. Furthermore, it lacks a rigorous mathematical foundation for its empirical mode decomposition (EMD), which undermines the stability and reliability of its instantaneous frequency representation.
@@
To address these limitations, we propose a framework that combines CWT-based time-frequency spectrograms with You Only Look Once (YOLO)-based object detection models, including YOLOv9 [11], YOLOv10 [12], and YOLOv11 [13] to localize and classify fault patterns simultaneously. This approach leverages YOLO’s unified detection pipeline, which is optimized for both accuracy and speed, to enables fast and spatially explainable bearing fault diagnostics. To the best of our knowledge, this is the first study to explore YOLO-based spatial detection in time-frequency spectrograms for industrial bearing fault applications. The main contributions of this letter are as follows:
  • We propose a novel bearing fault diagnosis framework that combines CWT with YOLOv9, v10, and v11 to detect and classify fault patterns directly on time-frequency spectrograms.
@@ II-A Continuous Wavelet Transform (CWT)
In this work, we adopt the Morlet wavelet due to its favorable time-frequency localization and robustness to high-frequency noise. The Morlet wavelet combines a sinusoidal carrier with a Gaussian envelope $\psi(t)=e^{-\frac{t^{2}}{2}}\cos(5t)$, which offers balanced resolution across different frequency bands and has shown superior performance in early fault detection [7]. By applying CWT to raw vibration signals, we generate two-dimensional time-frequency spectrograms that expose transient features and scale-dependent energy distributions. These spectrograms serve as spatially structured inputs for subsequent object detection models.
II-B Time-Frequency Spectrogram Visualization
To validate the visual enhancement provided by CWT, we convert the same vibration signals from Fig. 1 into two-dimensional time-frequency spectrograms using Morlet-based wavelet transformation. The resulting images, shown in Fig. 2, capture distinct energy concentration patterns that correspond to different fault types.
@@ II-C1 Signal Segmentation and Spectrogram Generation
For each dataset (CWRU, PU, and IMS), the raw vibration signals are segmented into fixed-length windows, each containing a sufficient number of samples to capture fault-relevant transients. Empirically, a window length of $2048$ samples with $50\%$ overlap is used to balance time resolution and data diversity. Each segment is transformed into a 2D time-frequency spectrogram using the Morlet-based CWT described earlier. The resulting grayscale spectrograms are logarithmically compressed and normalized to a $[0,1]$ range.
@@ II-C2 Bounding Box Annotation for Object Detection
Unlike traditional classification models that produce a single label for an entire signal segment, YOLO models require localized object annotations in the form of bounding boxes. To enable this, we manually annotate each spectrogram using the LabelImg tool, enclosing the time-frequency regions that exhibit fault-specific energy concentrations. Each box is associated with a class label corresponding to one of four fault types: Normal, Ball fault, Inner race fault, and Outer race fault.
@@ III-C Performance Comparison Across Datasets
Compared to YOLOv9, YOLOv11 reduces FLOPs by over 97% and parameters by nearly 95%, while achieving comparable or better accuracy across all datasets. Although MCNN-LSTM remains the most compact model in terms of complexity, its classification accuracy and generalizability are significantly lower. In contrast, YOLOv11 provides a balanced trade-off between efficiency and high detection accuracy, making it suitable for lightweight industrial fault monitoring. The proposed CWT-YOLO framework demonstrates strong generalization across diverse datasets. On the PU dataset, which includes multiple fault types under varying load and speed, YOLOv11 outperforms the baseline by over 16 percentage points in mAP. On the IMS dataset, all YOLO models achieve near-perfect classification, highlighting their ability to capture subtle degradation over long-term operation. Among the YOLO variants, YOLOv11 achieves the best overall performance, combining lower model complexity with faster inference than YOLOv9 and YOLOv10. These results indicate that the spatial detection paradigm, when applied to time-frequency spectrograms, supports accurate and interpretable fault localization.
@@ IV Conclusion
This letter presents a bearing fault diagnosis framework that combines CWT with YOLO-based object detection to leverage time-frequency spectrograms for localized classification. By converting one-dimensional vibration signals into spatially structured spectrograms, the system enables visual identification of fault patterns using bounding box detection. Evaluations on the CWRU, PU, and IMS datasets confirm that YOLO-based models significantly outperform the MCNN-LSTM baseline in classification accuracy, with YOLOv11 achieving the best balance between detection performance and computational complexity. These results highlight the potential of spectrogram-level object detection for building practical, interpretable, and deployment-ready condition systems.